Solve the inequality -5≤ -2x+ 2/3≤ 4
Answer by Boreal(15235)

-5 ≤ -2x+ 2/3 ≤ 4
Multiply everything by 3 to clear fractions
-15 <= -6x+2 <=12
-6x+2 < = 12
-6x < =10
x >= -5/3
-15 <= -6x+2
-17 <= -6x
(17/6) >= x
x < =(17/6)
[(-5/3),(17/6)]
check
-2 should not work
-5 <= 4+(4/3) < = 4 NOT TRUE
0 should work
-5 <=2/3 < = 4 TRUE
3 should not work
-5 <=-6+ 2/3 < = 4 NOT TRUE